# Index Terms: Driving automation systems, safety impact assessments, equivalence testing, synthetic test scenario validationI Introduction

### v1

@@ II Equivalence Testing
| Method procedure | Data distribution model | Test scope | Extend to the multivariate use case | |||
@@
| Frequentist TOST-based [16] | Parametric | Mean or/and variance | EDNE [17], SE [17] | [18, 19] | [20] | |
| Nonparametrica | Median (MWW [21, 22]) | Generalized MWW [23] | ||||
@@
| Bayesian | ROPE-based [15] | Parametric | Estimated model parameters (e.g., mean, variance, and regression coefficients) | EDNE [17], SE [17] | [18, 19] | [30] |
@@ II-2 Bayesian approaches
II-3 Extensions to the Multivariate Use Case
@@
The second category of methods is multiple comparison tests (MCTs) [18, 19]. In MCTs, the two data groups are declared equivalent only if all tested statistics are within the equivalence regions. When conducting multiple equivalence tests, the chance of making a false nonequivalence error increases if each test is carried out at the same stringent level as when only one test is performed. Therefore, adopting less restrictive thresholds for multiple individual tests is advisable.
@@ III-D Step 4: Overall equivalence assessment
Practical equivalence testing relies heavily on expert judgment and interpretive reasoning rather than purely numerical outcomes. Therefore, the rationale behind declaring equivalence or nonequivalence must be thoroughly documented. This documentation should clearly outline the motivations behind the selection of specific metrics, statistics, ROPEs, and the overall equivalence criteria.
@@ IV-B Metrics and statistics
Choosing appropriate metrics and statistics is crucial, since they must capture the scenario characteristics that are most relevant to the intended assessment. As noted, these choices should be guided by expert judgment after a thorough analysis of each characteristic’s impact on the performance of the system under assessment.
For demonstration purposes, we selected four metrics that capture crash severity, time available for evasive maneuvers, and vehicle braking behaviors:
[m/s]: The estimated lead vehicle’s total velocity change (Delta-v) over the crash event.
[s]: The no-return time, marking the point of no return beyond which a collision is unavoidable even if the following vehicle applies the maximum deceleration of -9 m/s2 [41]. Time zero is the impact moment; thus, has a negative value.
[m/s2]: The lead vehicle’s minimum acceleration (or maximum deceleration).
[m/s2]: The following vehicle’s minimum acceleration.
@@
The remaining two metrics and describe the braking behaviors of the leading and following vehicles, respectively. Fig. 2 shows the weighted cumulative density functions of all four metrics for the reference and PCM datasets. Unlike and , and each contain a point mass at 0 m/s2, which indicates two vehicle-behavior patterns (i.e., non-braking and braking). To measure the patterns separately, we used the following two statistics for each of the two metrics: 1) the ratio of proportions of non-braking (i.e., = 0 m/s2) scenarios (, similar to ) and 2) the KS statistic for braking (i.e., excluding nonbraking cases 0 m/s2) scenarios ().
@@ IV-C Equivalence criteria
It is important to note that the ROPEs chosen in this work are purely illustrative and are neither linked to a specific assessment context nor presented as recommendations. In practice, the ROPEs should be determined through expert judgment after a thorough analysis of each metric’s impact on the performance of the system under assessment; for instance, the ROPE’s stringency should increase with the relevance/importance of the corresponding statistic and metric. On the other hand, as noted, adopting less restrictive ROPEs for a metric with multiple statistics is advisable to mitigate the risk of false nonequivalence conclusions. In our example, we treated as the most critical metric, given its relevance to safety-impact assessments of driving automation systems, followed by as the second most important metric. As a result, we set narrower ROPEs for statistics corresponding to these two metrics (see Table II). For the two less important metrics ( and ), we chose larger ROPEs.
@@ IV-D Results
| Metric | Equivalence test | Significance test | |||||||
@@
For the equivalence tests, all the statistics passed the equivalence tests, indicating the practical equivalence of this metric. However, the results show nonequivalence of the other three metrics. Meanwhile, all significant tests except the one for demonstrated a statistically significant difference at the 0.05 significance level. The comparison of the equivalence and significance test results reveals no obvious connection: a statistically nonsignificant difference does not imply practical equivalence (as seen in the test results for ), while a statistically significant difference may still be practically negligible (as seen in the test results for ). In summary, given the results that only one metric (), a less-critical one, passed the equivalence test while the other metrics failed, the demonstration indicates that the two datasets are not considered practically equivalent.
@@ V Discussion and conclusions
Of note, several key practical considerations exist when conducting an equivalence test. First, it is crucial to select the right metrics: they must represent the scenario characteristics that are most relevant for the intended assessment. Using more metrics adds comparison dimensions, making the test more thorough while also increasing the chance of making a false nonequivalence error. Therefore, choosing only the most relevant metrics for the assessment scope is essential, as it helps avoid mistakenly declaring nonequivalence due to nonequivalent, less-relevant/important metrics.
@@
A key limitation of this study is that the metrics were modeled and evaluated independently, ignoring potential correlations between them. Additionally, our method relies on separate equivalence tests for individual metrics rather than employing a unified multivariate ROPE framework. The resulting conservative outcomes increase the risk of false nonequivalence conclusions. Future research should consider multivariate Bayesian ROPE methods that explicitly account for interdependent metrics, which could enhance the accuracy and robustness of practical equivalence testing.

### v3

# IEEE Copyright Notice

@@ II Equivalence Testing
Method procedure |
Data distribution model |
Test scope |
Extend to the multivariate use case | |||
@@
Frequentist TOST-based [16] |
Parametric | Mean or/and variance | EDNE [17], SE [17] |
[18, 19] |
[20] | |
Nonparametrica |
Median (MWW [21, 22]) | Generalized MWW [23] |
@@
Bayesian |
ROPE-based [15] | Parametric | Estimated model parameters (e.g., mean, variance, and regression coefficients) | EDNE [17], SE [17] | [18, 19] | [30] |
@@ II-B Bayesian Approaches
II-C Extensions to the Multivariate Use Case
@@
The second category of methods is multiple comparison tests (MCTs) [18, 19]. In MCTs, the two data groups are declared equivalent only if all tested statistics are within the equivalence regions. When conducting multiple equivalence tests, the chance of making a false non-equivalence error increases if each test is carried out at the same stringent level as when only one test is performed. Therefore, adopting less restrictive thresholds for multiple individual tests is advisable.
@@ III-D Step 4: Overall Equivalence Assessment
Practical equivalence testing relies heavily on expert judgment and interpretive reasoning rather than purely numerical outcomes. Therefore, the rationale behind declaring equivalence or non-equivalence must be thoroughly documented. This documentation should clearly outline the motivations behind the selection of specific metrics, statistics, ROPEs, and the overall equivalence criteria.
@@ IV-B Metrics and Statistics
Choosing appropriate metrics and statistics is crucial, since they must capture the scenario characteristics most relevant to the intended assessment. As noted, these choices should be guided by expert judgment after a thorough analysis of each characteristic’s impact on the performance of the system under assessment.
For demonstration purposes, we selected four general metrics that capture crash severity, time available for evasive maneuvers, and vehicle braking behaviors:
: The estimated lead vehicle’s total velocity change (Delta-v) over the crash event.
: The no-return time, marking the point of no return beyond which a collision is unavoidable even if the following vehicle applies the maximum deceleration of [41]. Time zero is the impact moment; thus, has a negative value.
: The lead vehicle’s minimum acceleration (or maximum deceleration).
: The following vehicle’s minimum acceleration.
@@
The remaining two metrics and describe the braking behaviors of the leading and following vehicles, respectively. Fig. 2 shows the weighted cumulative density functions of all four metrics for the reference and PCM datasets. Unlike and , and each contain a point mass at 0 m/s2, which indicates two vehicle-behavior patterns (i.e., non-braking and braking). To measure the patterns separately, we used the following two statistics for each of the two metrics: 1) the ratio of proportions of non-braking (i.e., = 0 m/s2) scenarios (, similar to ) and 2) the KS statistic for braking (i.e., excluding non-braking cases 0 m/s2) scenarios ().
@@ IV-C Equivalence Criteria
It is important to note that the ROPEs chosen in this work are purely illustrative and are neither linked to a specific assessment context nor presented as recommendations. In practice, the ROPEs should be determined through expert judgment after a thorough analysis of each metric’s impact on the performance of the system under assessment; for instance, the ROPE’s stringency should increase with the relevance/importance of the corresponding statistic and metric. On the other hand, as noted, adopting less restrictive ROPEs for a metric with multiple statistics is advisable to mitigate the risk of false non-equivalence conclusions. In our example, we treated as the most critical metric, given its relevance to safety-impact assessments of driving automation systems, followed by as the second most important metric. As a result, we set narrower ROPEs for statistics corresponding to these two metrics (see Table II). For the two less important metrics ( and ), we chose larger ROPEs.
@@ IV-D Results
Metric |
Equivalence test | Significance test | |||||||
@@
For the equivalence tests, all the statistics passed the equivalence tests, indicating the practical equivalence of this metric. However, the results show non-equivalence of the other three metrics. Meanwhile, all significant tests except the one for demonstrated a statistically significant difference at the 0.05 significance level. The comparison of the equivalence and significance test results reveals no obvious connection: a statistically nonsignificant difference does not imply practical equivalence (as seen in the test results for ), while a statistically significant difference may still be practically negligible (as seen in the test results for ). In summary, given the results that only one metric (), a less-critical one, passed the equivalence test while the other metrics failed, the demonstration indicates that the two datasets are not considered practically equivalent.
@@ V Discussion and conclusions
Of note, several key practical considerations exist when conducting an equivalence test. First, it is crucial to select the right metrics: they must represent the scenario characteristics that are most relevant for the intended assessment. Using more metrics adds comparison dimensions, making the test more thorough while also increasing the chance of making a false non-equivalence error. Therefore, choosing only the most relevant metrics for the assessment scope is essential, as it helps avoid mistakenly declaring non-equivalence due to non-equivalent, less-relevant/important metrics.
@@
A key limitation of this study is that the metrics were modeled and evaluated independently, ignoring potential correlations between them. Real-world crash scenario metrics, such as crash severity () and temporal criticality (), often exhibit inherent correlations due to their shared underlying dynamics. Treating these metrics independently could result in overly conservative practical equivalence decisions and a higher incidence of false non-equivalence conclusions. Future work should address this limitation by employing multivariate Bayesian ROPE-based equivalence testing methods. Such methods explicitly consider the joint distributions and interactions among metrics, enhancing the accuracy and reliability of equivalence assessments.
In addition, the sensitivity of the ROPE method to prior assumptions is important, especially in regulatory contexts where consistency and transparency are essential. Future studies should conduct explicit sensitivity analyses to quantify how the choice of priors affects equivalence conclusions.
Despite these limitations, the proposed method is an important initial step toward a systematic and transparent practical equivalence test between synthetic and real-world pre-crash datasets. In fact, the method provides practitioners with a systematic and rigorous tool to validate synthetic scenario generation processes by clearly quantifying similarities and discrepancies between synthetic and real-world datasets. It enables targeted identification of aspects needing refinement, thus guiding iterative improvements in scenario realism and representativeness.
Acknowledgment
This research was funded by FFI Vinnova, a Swedish governmental agency for innovation, as part of the project Improved quantitative driver behavior models and safety assessment methods for ADAS and AD (QUADRIS: nr. 2020-05156).